# Carroll black holes in (A)dS and their higher-derivative modifications

### v2

@@ I Introduction
This theory has the Schwarzschild metric as a solution; however, it also contains a one-parameter family of static spherically symmetric solutions called the Schwarzschild-Bach black holes whose deviation from Schwarzschild is measured by a quantity called the Bach parameter denoted by $\delta$ [8, 9, 10, 11]. These black holes are static spherically symmetric solutions of the field equations, thus, their metric is given by the form
@@
where $\omega$ and $t$ are parameters in the Frobenius series that will be used to differentiate different classes of solutions. It was shown in [9] that there are three classes of Frobenius solutions: The first is defined by setting $\omega=t=0$, called class $(0,0)$, which has no horizon, the second is defined by $\omega=1$ and $t=0$, called class $(1,0)$, which defines a wormhole, and the third is defined by $\omega=t=1$, called class $(1,1)$, which contains the Schwarzschild solution as well as the Schwarzschild-Bach solutions; for more details on such solutions, we refer the reader to [12]. There are also other classes derived from non-Frobenius series, but we will not discuss those. In this paper, we are only interested in the class $(1,1)$ since it contains the class of Schwarzschild-Bach black holes. In this work, we will construct the Carrollian versions of these black holes.
@@
The Carrollian limit is defined as the ultra-local limit of the Poincaré group where the speed of light $c$ goes to zero, $c\to 0$. It was defined independently by Levy-Leblond [13] and Sen Gupta [14] in the 1960s as the Inönü-Wigner contraction of the Poincaré group. There was not much work in this direction until recently due to [15], where it was shown to be related to near-horizon physics. Since then, it has been studied from various perspectives and applied in various areas of physics [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 28, 51, 52, 53, 54].
@@
Here, we will focus on Carroll black holes, which were first properly introduced in [55] and further analyzed in [56]. They are defined on a Carrollian manifold, which is a manifold endowed with a nowhere vanishing vector field $v^{\mu}$ and a degenerate metric $h_{\mu\nu}$ such that $v^{\mu}h_{\mu\nu}=0$. Furthermore, we define an associated form $E_{\mu}$ to $v^{\mu}$ such that $E_{\mu}v^{\mu}=1$. Notice that $E_{\mu}$ plays the role of an Ehresmann connection and need not be covariantly constant. A Carrollian manifold can be defined using a Lorentzian manifold $\mathcal{M}$ by splitting its tangent bundle $T\mathcal{M}$ into a vertical bundle $\mathrm{Ver}\mathcal{M}$ (a vector bundle whose fibers are isomorphic to the time coordinate) and horizontal bundles $\mathrm{Hor}\mathcal{M}$ (whose fibers are isomorphic to metrics on spacelike slices of $\mathcal{M}$), i.e, $T\mathcal{M}=\mathrm{Ver}\mathcal{M}\oplus\mathrm{Hor}\mathcal{M}$. This results in a foliation of the manifold into a codimension 1 submanifold with metric $h_{\mu\nu}$ as leaves. We denote the adapted coordinates in the leaves by $x_{\perp}^{\mu}$, and the coordinate in direction of $v^{\mu}$ by $x^{\mu}_{\parallel}$. A detailed discussion of Carrollian geometry and Carrollian symmetries can be found in [57, 58, 59]. Carrollian manifolds can be endowed with differential structures that require both $v^{\mu}$ and $h_{\mu\nu}$ to be covariantly constant. This defines special torsional connections; see [44] for more details. A Carroll black hole defined as a solution of a Carroll gravity theory (a gravity theory defined on a Carroll manifold) that has a Carroll extremal surface (a surface for which the norm of $v^{\mu}$ is zero) and has certain thermodynamical properties stated in [55].
@@
In this paper, we introduce the Carrollian version of Schwarzschild-(A)dS and Schwarzschild-Bach-(A)dS black holes as solutions to the Carroll limit of GR and quadratic gravity and analyze the geodesic motion on their geometries. We show that, in the case of Carroll Schwarzschild-(A)dS, a near-tangential particle winds around the extremal surface with a finite number of windings, similar to the Carrol Schwarzschild black hole, but this number receives corrections due to the cosmological constant. On the other hand, in the case of Schwarzschild-Bach-(A)dS black hole, the number of windings is infinite regardless of the cosmological constant, i.e. this feature is only due to the higher-curvature terms added to the Lagrangian. We also study the thermodynamical properties of such black holes and compare them to the incompressible thermodynamical system with a divergent entropy and vanishing temperature. In what follows, we introduce a scale for each dimensionful quantity and work with the dimensionless quantities obtained by dividing by these scales. The paper is organized as follows:
@@ II.1 Geodesics
We start by computing the Carroll limit of Schwarzschild-Bach black hole and then derive the geodesic equations for a test particle in that geometry following the procedure in [56]. The metric for the Schwarzschild-Bach black hole is (3). We choose the following foliation:
$h_{\mu\nu}dx^{\mu}dx^{\nu}=\tfrac{dr^{2}}{f(r)}+r^{2}d\Omega^{2},$ $\ \ \ E_{\mu}dx^{\mu}=-h(r)dt,$ $v^{\mu}\partial_{\mu}=\tfrac{-1}{h(r)}\partial_{t}.$ (5)
One can easily see that $h_{\mu\nu}v^{\mu}=0$ and $E_{\mu}v^{\mu}=1$ which demonstrate the fact that it is indeed a Carrollian construction. By straightforward calculations, we can also verify that (5) satisfies the Hamiltonian constraints and the magnetic equations of motion of the theory labeled (2,4) in [60]. That is a Carrollian theory derived from the limit of quadratic gravity where the parameters of the action (1) are assumed to be of the form $\alpha=c^{2}\alpha^{\prime}$ and $\beta=c^{4}\beta^{\prime}$, where $\alpha^{\prime}$ and $\beta^{\prime}$ have no $c$ dependencies. This means that the chosen foliation gives a solution for the Carrollian theory, and since it has a Carroll extremal surface corresponding to $f(r)=0$ (where the degenerate metric diverges, giving rise to a structural singularity [55]), this solution can be called a Carroll black hole whose metric is defined near the extremal surface. Thus, we expand $f(r)$ into powers of $r-r_{0}$ as follows:
@@
[…]
where $\dot{x}=\tfrac{dx}{d\tau}$. In this paper, we choose $x^{\mu}_{\perp}=r,\theta,\phi$ to be the spherical coordinates on the spatial slice, $x^{\mu}_{\parallel}=t$ is the time coordinate (normal to the spatial slice), and $g_{0},g_{1},g_{2}$ are arbitrary constants. As we will see below, this action matches the Carrollian limit of the action of geodesic particles in the Lorentzian case. Without loss of generality, we assume that the motion is in the equatorial plane $\theta=\tfrac{\pi}{2}$. Substituting (5) into (7), we get
[…]
@@
Using the above equations, we can now derive the constraint for this action to match the Carroll limit of the Lorentzian geodesic action by varying the action with respect to $e$, the result is
@@
[…]
which matches the Carroll limit of the relativistic geodesic particle (up to an overall constant) with different coupling constants for the electric and the magnetic parts. Following [56], the definition of the particle’s energy is chosen to be
$E=\tfrac{g_{0}}{2g_{1}}e^{2}-\tfrac{g_{2}}{2g_{1}}F^{2},$ (14)
which, apart from the coefficients of $e^{2}$ and $F^{2}$, can be derived from the Carroll limit of the Lorentzian energy for a geodesic particle. The presence of the coefficients indicates that not all Carrollian formulae can be derived as a limit of a Lorentzian one. Using this definition, the geodesic equation is written as
@@
where $V^{\textrm{eff}}(r)=\tfrac{l^{2}f(r)}{2r^{2}}-(f(r)-1)E$ is the effective potential. Notice that when $f(r)=1-\tfrac{2M}{r}$, we recover the results for Schwarzschild derived in [56].
Now, we write the circularity conditions as
@@
×$\times F\bigg{(}\sin^{-1}\bigg{(}\frac{\sqrt{\frac{b+\frac{1}{u}}{b}}}{\sqrt{2}}\bigg{)}|\frac{2bu_{0}}{bu+1}\bigg{)}\bigg{)}\bigg{)}.$
where $\Pi(n;x|m)$ is the incomplete elliptic integral of the third kind and $F(x|n)$ is the elliptic integral of the first kind. We should also take the limit $b\to 1/u_{0}$ to ensure that the particle enters the area where the solution is valid. To obtain the deflection angle of the particle, we evaluate this function at a point and take the difference to the value of the function at the turning point $u=1/b$. The problem with this procedure is that (27) diverges at $u=1/b$. In an attempt to get a finite value from the above formula, we introduce a regularization parameter $a$ so that the function is evaluated at $1/b-a$. The result diverges in the limit $a\to 0$ for any choice of the parameters in the theory, namely $k$ and $\delta$. This is mainly due to the divergence of $F(x|n)$ and $\Pi(n;x|n)$ in the limits $x\to 1$ and $n\to 1$, which are precisely the limits $a\to 0$ and $b\to u_{0}$, necessary for the consistency of the calculations. Thus, we can conclude that the particle undergoes an infinite number of windings (since the deflection angle is infinite) as shown in Fig. 1, i.e. the addition of higher derivative terms to the action results in the situation where particles are trapped near the extremal surface if they come close enough. This can be compared to Carroll Schwarzschild’s case where a particle winds a finite number of times and then gets ejected to asymptotic infinity as shown in Fig. 1.
@@ II.2 Thermodynamics
$S_{(2)}=-\int d^{4}x$ $e\big{\{}-\bar{R}+\beta^{\prime}\big{[}(K^{2}-K_{\mu\nu}K^{\mu\nu})(K^{2}-K_{\mu\nu}K^{\mu\nu}+4\pounds_{\boldsymbol{v}}K)-4(\pounds_{\boldsymbol{v}}K)^{2}\big{]}+\alpha^{\prime}\big{[}-2\bar{R}{}^{\lambda\nu}\nabla_{\mu}(V^{\mu}K_{\lambda\nu})$ (33)
@@
$S_{(3)}=\int d^{4}xe$ $\bigg{(}\tfrac{1}{4}B^{\mu\nu}B_{\mu\nu}-\alpha^{\prime}\big{(}\tfrac{1}{2}\mathcal{K}^{\alpha\beta}\mathcal{K}_{\alpha\beta}B^{\mu\nu}B_{\mu\nu}-\Pi^{\nu\alpha}\nabla_{\alpha}\mathcal{K}\nabla_{\rho}(B_{\nu}{}^{\rho})-\tfrac{1}{4}V^{\sigma}\mathcal{K}^{\nu\rho}B_{\sigma\rho}\nabla_{\mu}(B_{\nu}{}^{\mu})$ (36)
@@ III.1 Geodesics
[…]
@@
[…]
The first integral is the same as that of Schwarzschild and the second integral will serve as a correction due to the presence of the cosmological constant. The second integral diverges in the upper limit, so we introduce a regularization parameter $a$, integrate from 0 to $1/b-a$, then expand as a series of powers of $a$, keeping only the leading term. We will see that by appropriately fixing $a$, we get a finite deflection angle. The result of the integrals is

### v3

@@ I Introduction
This theory has the Schwarzschild metric as a solution; however, it also contains a one-parameter family of static spherically symmetric solutions called the Schwarzschild-Bach black holes whose deviation from Schwarzschild is measured by a quantity called the Bach parameter denoted by $\delta$ [8, 9, 10, 11]. Such black holes are static spherically symmetric solutions of the field equations. Thus, their metric is given by the form
@@
where $\omega$ and $t$ are parameters in the Frobenius series that will be used to differentiate different classes of solutions. As shown in [9], there are three classes of Frobenius solutions:
  1. 1.
    The class defined by setting $\omega=t=0$, denoted by $(0,0)$, which has no horizon.
  2. 2.
    The class defined by setting $\omega=1$ and $t=0$, denoted by $(1,0)$, which defines a wormhole.
  3. 3.
    the The class defined by setting $\omega=t=1$, denoted by $(1,1)$, which contains the Schwarzschild solution as well as the Schwarzschild-Bach solutions.
@@
For more details on the aforementioned classes and solutions, we refer the reader to [12]. There are also other classes derived from non-Frobenius series, but we will not discuss those. In this paper, we are only interested in the class $(1,1)$ since it contains the class of Schwarzschild-Bach black holes. In this work, we will construct the Carrollian versions of these black holes.
@@
The Carroll group is defined as the ultra-local limit of the Poincaré group where the speed of light $c$ goes to zero, $c\to 0$. It was defined independently by Levy-Leblond [13] and Sen Gupta [14] in the 1960s as the Inönü-Wigner contraction of the Poincaré group. Afterwards, it was recognised that this group appears in near black hole horizon physics [15, 16, 17] and at null infinity [18, 19, 20, 21, 22]. Since then, it has been studied from various perspectives and applied in various areas of physics [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 35, 58, 59, 60, 61, 62].
@@
Despite all the work in this direction Carroll black holes, which were first properly introduced in [63] were rarely studied (for example in [64]). They are defined on a Carrollian manifold, which is a manifold endowed with a nowhere vanishing vector field $v^{\mu}$ and a degenerate metric $h_{\mu\nu}$ such that $v^{\mu}h_{\mu\nu}=0$. Furthermore, we define an associated form $E_{\mu}$ to $v^{\mu}$ such that $E_{\mu}v^{\mu}=1$. Notice that $E_{\mu}$ plays the role of an Ehresmann connection and need not be covariantly constant. A Carrollian manifold can alternatively be defined by contracting the light cone at each point of a Lorentzian manifold $\mathcal{M}$. This splits its tangent bundle $T\mathcal{M}$ into a vertical bundle $\mathrm{Ver}\mathcal{M}$ (a vector bundle whose fibers are isomorphic to the timelike vector flow) and horizontal bundles $\mathrm{Hor}\mathcal{M}$ (whose fibers are isomorphic to spacelike slices of $\mathcal{M}$), i.e, $T\mathcal{M}=\mathrm{Ver}\mathcal{M}\oplus\mathrm{Hor}\mathcal{M}$. This results in a foliation of the manifold, by lines corresponding to the flow of the kernel of the induced metric i.e. $v^{\mu}$, into codimension 1 submanifolds with metric $h_{\mu\nu}$ as leaves (as reminiscents of the spacelike slices in the original Lorentzian manifold). We note that is different from the foliation of Lorentzian manifolds into codimension 1 submanifolds without the light cone contraction since the timelike vector flow is not restricted from one leave to the other. We denote the adapted coordinates in the leaves by $x_{\perp}^{\mu}$, and the coordinate in direction of $v^{\mu}$ by $x^{\mu}_{\parallel}$. A detailed discussion of Carrollian geometry and Carrollian symmetries can be found in [65, 66, 67]. Carrollian manifolds can be endowed with differential structures that require both $v^{\mu}$ and $h_{\mu\nu}$ to be covariantly constant. This defines special torsional connections; see [51] for more details. A Carroll black hole defined as a solution of a Carroll gravity theory (a gravity theory defined on a Carroll manifold) that has a Carroll extremal surface (a surface for which the norm of $v^{\mu}$ is zero) and has certain thermodynamical properties stated in [63].
@@
In this paper, we study Carroll black holes further as geometries and as thermodynamical systems. We introduce the Carrollian version of Schwarzschild-(A)dS and Schwarzschild-Bach-(A)dS black holes as solutions to the Carroll limit of GR and quadratic gravity and analyze the geodesic motion on their geometries. We show that, in the case of Carroll Schwarzschild-(A)dS, a near-tangential particle winds around the extremal surface with a finite number of windings, similar to the Carroll Schwarzschild black hole, but this number receives corrections due to the cosmological constant. On the other hand, in the case of Schwarzschild-Bach-(A)dS black hole, the number of windings is infinite regardless of the cosmological constant, i.e. this feature is only due to the higher-curvature terms added to the Lagrangian. We also study the thermodynamical properties of such black holes and compare them to the incompressible thermodynamical system with a divergent entropy and vanishing temperature. In what follows, we introduce a scale for each dimensionful quantity and work with the dimensionless quantities obtained by dividing by these scales. The paper is organized as follows:
@@ II.1 Geodesics
We start by computing the Carroll limit of Schwarzschild-Bach black hole and then derive the geodesic equations for a test particle in that geometry following the procedure in [64]. The metric for the Schwarzschild-Bach black hole is (3), its Carroll limit is derived by restoring the factors of $c$ so it can be written as
$ds^{2}=-h(r)c^{2}dt^{2}+\tfrac{dr^{2}}{f(r)}+r^{2}d\Omega^{2},$ (5)
then taking the limit $c\to 0$. The result is a degenerate metric, and a vector in its kernel (along with its associated one form) given by
$h_{\mu\nu}dx^{\mu}dx^{\nu}=\tfrac{dr^{2}}{f(r)}+r^{2}d\Omega^{2},$ $\ \ \ v^{\mu}\partial_{\mu}=\tfrac{-1}{h(r)}\partial_{t},$ $E_{\mu}dx^{\mu}=-h(r)dt$ (6)
Notice that due to the fact that the Lorentzian metric is static, the Carroll limit looks like a partition of the Lorentzian data. However, one can easily check that $h_{\mu\nu}v^{\mu}=0$ and $E_{\mu}v^{\mu}=1$ which demonstrate the fact that it is indeed a Carrollian construction rather than just a Lorentzian data partition. By straightforward calculations, we can also verify that (6) satisfies the Hamiltonian constraints and the magnetic equations of motion of the theory labeled (2,4) in [68]. That is a Carrollian theory derived from the limit of quadratic gravity where the parameters of the action (1) are assumed to be of the form $\alpha=c^{2}\alpha^{\prime}$ and $\beta=c^{4}\beta^{\prime}$, where $\alpha^{\prime}$ and $\beta^{\prime}$ have no $c$ dependencies. This means that the chosen foliation gives a solution for the Carrollian theory, and since it has a Carroll extremal surface corresponding to $f(r)=0$ (where the degenerate metric diverges, giving rise to a structural singularity [63]), this solution can be called a Carroll black hole whose metric is defined near the extremal surface. Thus, we expand $f(r)$ into powers of $r-r_{0}$ as follows:
@@
$I=\int d\tau e(g_{0}+g_{1}e^{-2}h_{\mu\nu}\dot{x}^{\mu}_{\perp}\dot{x}_{\perp}^{\nu}+g_{2}e^{-2}\dot{x}^{2}_{\parallel}),$ (8)
where $\dot{x}=\tfrac{dx}{d\tau}$. This action is Carroll diffeomorphism invariant since only Carroll diffeomorphism invariant quantities couple to $\dot{x}^{\mu}$. Notice that we do not require local Carroll boost symmetry which forbids local energy flux and in turn forbids the motion of massive particles. In our case such symmetry is broken so we expect non trivial dynamics. In this paper, we choose $x^{\mu}_{\perp}=r,\theta,\phi$ to be the spherical coordinates on the spatial slice, $x^{\mu}_{\parallel}=t$ is the time coordinate (normal to the spatial slice), and $g_{0},g_{1},g_{2}$ are arbitrary constants. Without loss of generality, we assume that the motion is in the equatorial plane $\theta=\tfrac{\pi}{2}$. Substituting (6) into (8), we get
$I=\int d\tau e(g_{0}+g_{1}e^{-2}(\tfrac{\dot{r}^{2}}{f(r)}+r^{2}\dot{\phi}^{2})+g_{2}e^{-2}\dot{t}^{2}).$ (9)
@@
Using the above equations, we can now derive the constraint for this action due to reparametrization invariance by varying the action with respect to $e$, the result is
@@
$I=\int 2\sqrt{g_{1}(\tfrac{\dot{r}^{2}}{f(r)}+l^{2})+g_{2}F^{2}}d\tau.$ (14)
Following [64], the definition of the particle’s energy is chosen to be
$E=\tfrac{g_{0}}{2g_{1}}e^{2}-\tfrac{g_{2}}{2g_{1}}F^{2}.$ (15)
The presence of the coefficients in the previous formulae indicates that not all Carrollian formulae can be derived as a limit of a Lorentzian one. Indeed, the formulae derived from the Carroll limit of the corresponding Lorentzian formulae must have fixed ratios of coefficients rather than arbitrary, namely, $\tfrac{g_{1}}{g_{2}}=\mathcal{O}(c^{-2})$ which implies that one can only have only the electric or the magnetic terms in the action, and the following formulae as proven in [65]. Here, we consider Carroll diffeomorphism actions without referring to a parent Lorentzian manifold. This allows us to add Carroll diffeomorphism invariant terms with arbitrary coupling constants without worrying about the Lorentz invariance of any Lorentzian manifold. Hence having interactions between the electric and the magnetic terms in a consistent way giving rise to more interesting dynamics as we will see below.
Using this definition, the geodesic equation is written as
@@
where $V^{\textrm{eff}}(r)=\tfrac{l^{2}f(r)}{2r^{2}}-(f(r)-1)E$ is the effective potential. Notice that when $f(r)=1-\tfrac{2M}{r}$, we recover the results for Schwarzschild derived in [64]. Now, we write the circularity conditions as
@@
$\times F\bigg(\sin^{-1}\bigg(\frac{\sqrt{\frac{b+\frac{1}{u}}{b}}}{\sqrt{2}}\bigg)|\frac{2bu_{0}}{bu+1}\bigg)\bigg)\bigg).$
where $\Pi(n;x|m)$ is the incomplete elliptic integral of the third kind and $F(x|n)$ is the elliptic integral of the first kind. We should also take the limit $b\to 1/u_{0}$ to ensure that the particle enters the area where the solution is valid. To obtain the deflection angle of the particle, we evaluate this function at a point and take the difference to the value of the function at the turning point $u=1/b$. The problem with this procedure is that (28) diverges at $u=1/b$. In an attempt to get a finite value from the above formula, we introduce a regularization parameter $a$ so that the function is evaluated at 1/ba1/b-a. The result diverges in the limit $a\to 0$ for any choice of the parameters in the theory, namely $k$ and $\delta$. This is mainly due to the divergence of $F(x|n)$ and $\Pi(n;x|n)$ in the limits $x\to 1$ and $n\to 1$, which are precisely the limits $a\to 0$ and $b\to u_{0}$, necessary for the consistency of the calculations. Thus, we can conclude that the particle undergoes an infinite number of windings (since the deflection angle is infinite) as shown in Fig. 1, i.e. the addition of higher derivative terms to the action results in the situation where particles are trapped near the extremal surface if they come close enough. This can be compared to Carroll Schwarzschild’s case where a particle winds a finite number of times and then gets ejected to asymptotic infinity as shown in Fig. 1.
@@ II.2 Thermodynamics
$S_{(2)}=-\int d^{4}x$ $e\big\{-\bar{R}+\beta^{\prime}\big[(K^{2}-K_{\alpha\beta}K^{\alpha\beta})(K^{2}-K_{\mu\nu}K^{\mu\nu}+4\pounds_{\boldsymbol{v}}K)-4(\pounds_{\boldsymbol{v}}K)^{2}\big]+\alpha^{\prime}\big[-2\bar{R}{}^{\lambda\nu}\nabla_{\mu}(V^{\mu}K_{\lambda\nu})$ (34)
@@
$-\tfrac{1}{2}V^{\lambda}V^{\alpha}\nabla_{\mu}(B_{\alpha}{}^{\mu})\nabla_{\nu}(B_{\lambda}{}^{\nu})+\tfrac{3}{2}\bar{R}{}^{\alpha\lambda}\mathcal{K}^{\beta}_{\lambda}B_{\alpha\beta}+\bar{R}^{\mu\nu}\bar{R}_{\mu\nu}+\tfrac{1}{4}\mathcal{K}_{\beta\lambda}\mathcal{K}^{\rho\lambda}B^{\nu\beta}B_{\nu\rho}+\tfrac{1}{4}\mathcal{K}^{\beta\lambda}\mathcal{K}^{\alpha\rho}B_{\alpha\beta}B_{\lambda\rho}\big)$
@@ III.1 Geodesics
$\phi=-\int_{0}^{1/b}\frac{b}{\sqrt{1-b^{2}u^{2}}}\frac{1}{\sqrt{1-2Mu-\frac{\Lambda}{3u^{2}}}}du.$ (51)
@@
$\phi=-2\int_{0}^{1/b}\frac{b}{\sqrt{1-b^{2}u^{2}}}\frac{1}{\sqrt{1-2Mu}}du+2\Lambda\int_{0}^{1/b}\frac{\partial\psi}{\partial\Lambda}|_{\Lambda=0}\ du.$ (53)
The first integral is the same as that of Schwarzschild and the second integral will serve as a correction due to the presence of the cosmological constant. The second integral diverges in the upper limit, so we introduce a regularization parameter $a$, integrate from 0 to 1/ba1/b-a, then expand as a series of powers of $a$, keeping only the leading term. We will see that by appropriately fixing $a$, we get a finite deflection angle. The result of the integrals is